OpenStudy (lukecrayonz):

Why does this suddenly go away? Differentiate the function y=(8x-7)^4 I am using the power rule g(x)=8x-7 k=4 In the example, after some work it shows dy/dx=4(8x-7)^3*d/dx(8x-7) then =4(8x-7)^3*8 What happened?

OpenStudy (lukecrayonz):

@Psymon

OpenStudy (lukecrayonz):

OH, it's the deriv of 8x-7

OpenStudy (psymon):

Yeah, chain rule. YOu take the derivative of whats inside, too.

OpenStudy (lukecrayonz):

So dy/dx=32(8x-7)^3

OpenStudy (lukecrayonz):

So now for a new equation: y=(3x+4)^5

OpenStudy (lukecrayonz):

g(x)=3x+4 k=5

OpenStudy (lukecrayonz):

dy/dx=5(3x+4)^5-1*d/dx(3x+4)

OpenStudy (psymon):

Still going to be chain rule. You can kind of think as all derivatives as chain rule. You will always take the derivative of each inner function. If you have 3 exponents or things that barrier off expressions, then you have 3 derivatives you have to compute. Its just when you first do derivatives, the chain rule doesnt show up because the derivative of aninner function might be one and not matter. But yeah, keep going, lol.

OpenStudy (lukecrayonz):

d/dx(3x+4)=3

OpenStudy (lukecrayonz):

dy/dx=5(3x+4)^4*3 15(3x+4)^4

OpenStudy (psymon):

Looks good tome.

OpenStudy (lukecrayonz):

Hmm okay, so for a square root problem I will simply change the square root over to a power?

OpenStudy (psymon):

Yeah, realize a sqrt is a (1/2) power, meaning you will be multiplying by 1/2 and lowering the power to a -1/2

OpenStudy (lukecrayonz):

y=sqrt(8-7x) y=(8-7x)^1/2 k=0.5 g(x)=8-7x 0.5(8-7x)^-0.5*d/dx(8-7x) d/dx=8 0.5*8=4 4(8-7x)^-0.5

OpenStudy (lukecrayonz):

Hmm well mymathlab says it's wrong..

OpenStudy (psymon):

\[\sqrt{8-7x} = \frac{ 1 }{ 2 }(8-7x)^{-1/2}(-7)\] \[\frac{ -7 }{ 2\sqrt{8-7x} }\]

OpenStudy (lukecrayonz):

Ookay the example is way different :O? Weird.

OpenStudy (lukecrayonz):

Your answer is right, but how did you get -7 for d/dx?

OpenStudy (psymon):

derivative of 8-7x. Looks like you did it as if it were 8x-7 and not 8-7x.

OpenStudy (lukecrayonz):

Oh! Yes, I did. I was like how in the world did he get his constant into play?!

OpenStudy (psymon):

Lol, whoops xD

OpenStudy (lukecrayonz):

Okay:) y=(2x^2-5)^-10 k=-10 g(x)=2x^2-5 -10(2x^2-5)^-11*d/dx(2x^2-5) d/dx of 2x^2-5 drop the constant, 2x^2, bring down power, multiply, subtract power by 1 and you get 4x. so now, -10(2x^2-5)^-11*4x

OpenStudy (psymon):

Looks fine :3

OpenStudy (lukecrayonz):

Yikes! A factor! y=1/(2x-1)^3

OpenStudy (lukecrayonz):

Just help with the first step and then I can do it.. :)

OpenStudy (psymon):

Not bad, theres a trick to simply fraction ones: \[\frac{ -6 }{ (2x-1)^{4} }\] Just if you can see how I got that xD First step would be make the whole denominator into a negative exponent and put it in the numerator.

OpenStudy (lukecrayonz):

wat

OpenStudy (psymon):

Here's the first step: \[(2x-1)^{-3}\]

OpenStudy (lukecrayonz):

How did you get that?

OpenStudy (lukecrayonz):

Oh wait nevermind lo.

OpenStudy (psymon):

You can always change the sign of an exponent by moving it to the opposite side of a fraction. Lol, okay, haha.

OpenStudy (lukecrayonz):

I thought the original equation was (2x-1)^4 for a second, I was like What in the world did he just do!?

OpenStudy (psymon):

Theres a trick to simple fractions where I can give you derivatives of them very fast.

OpenStudy (lukecrayonz):

I'd like to learn it after this problem:)

OpenStudy (psymon):

If it makes sense, sure, lol.

OpenStudy (psymon):

Its not really a shortcut as much as it is pattern recognition and knowing whats going to happen.

OpenStudy (lukecrayonz):

Okay so after (2x-1)^-3, you do what?

OpenStudy (lukecrayonz):

Just chain rule it up and solve?

OpenStudy (psymon):

Do what youve been doing. k = -3, it multiplies down, lower power by 1, chain rule it out, etc xD

OpenStudy (lukecrayonz):

f(x)=(7+x^5)^5-(5+x^7)^3

OpenStudy (lukecrayonz):

Chain rule by themselves then subtract in the end?

OpenStudy (psymon):

Nothing new here. Yeah, work with each term separately,

OpenStudy (lukecrayonz):

(7+x^5)^5 k=5 g(x)=7+x^5 deriv: 5x^4 5(7+x^5)^4*5x^4 k=3 g(x)=5+x^7 deriv 7x^6 3(5+x^7)^2*7x^6 5(7+x^5)^4*5x^4-3(5+x^7)^2*7x^6

OpenStudy (psymon):

Yeah, looking good :3

OpenStudy (lukecrayonz):

Now how would you simplify this? Do I literally have to expand it or..

OpenStudy (psymon):

\[25x^{4}(7+x^{5})^{4}- 21x^{6}(5+x^{7})^{2}\]Yeah, hard to do much with it.

OpenStudy (lukecrayonz):

Hmm how did you get 25?

OpenStudy (psymon):

I have the same thing you have, its just multiplied a bit.

OpenStudy (lukecrayonz):

How old are you? Jw

OpenStudy (psymon):

Eh. 23.

OpenStudy (lukecrayonz):

You're very smart haha. f(x)=sqrt(x)+(x-3)^4

OpenStudy (psymon):

Sometimes yes, sometimes no, lol.

OpenStudy (psymon):

Same thing, just take the derivative of each termalone.

OpenStudy (lukecrayonz):

deriv of sqrt(x) is 1/2sqrt(x) So moving onto the next. 4*(x-3)^3*1 1/2sqrt(x)+4(x-3)^3

OpenStudy (psymon):

Right. Assuming you mean for the sqrt to be in the denominator, then yes

OpenStudy (lukecrayonz):

I'm only seventeen:( Haha

OpenStudy (psymon):

Well, Ive had more time to learn more? xD

OpenStudy (lukecrayonz):

y=x^2sqrt(10x-3)

OpenStudy (psymon):

SO finally a product rule, eh.

OpenStudy (lukecrayonz):

I know :P I just like to have an excuse for being dumb.

OpenStudy (lukecrayonz):

Alright so looking at the example this does not look that hard! f(x)=x^2 g(x)=sqrt(10x-3)

OpenStudy (psymon):

Haha, thats nice to have xD

OpenStudy (psymon):

Yeah, its just extra steps. Easy stuff, though.

OpenStudy (lukecrayonz):

Do you have a degree in math or a math-related field? And can you walk me through this? The problem has an error after the third step and won't load

OpenStudy (psymon):

Im working on a math degree, but im still a student xD Um...yeah, f = x^2 g = sqrt(10x-3) So derivative of f times regular g + derivative of g times regular f. A lot of text books show the reverse order, but both work. As far as Im concerned, f(x) is first, so Im doing f'(x) first, haha. SO yeah, full formula is f'(x)g(x) +f(x)g'(x) So doing that I get: \[2x \sqrt{10x-3}+x^{2}(10)(\frac{ 1 }{ 2 })(10x-3)^{-1/2}\]

OpenStudy (lukecrayonz):

That answer haha, when I was typing it in I was like "No way is this right, it's all over the place" Guess not ;_;

OpenStudy (psymon):

Nah, some of these get ridiculous, its normal xD

OpenStudy (lukecrayonz):

f(x)=((x-6)/(x+3))^5

OpenStudy (psymon):

Everything to the 5th, correct?

OpenStudy (lukecrayonz):

So yeah, it's a fraction of x-6 over x+3, all to the power of 5

OpenStudy (psymon):

Alrighty. Nothing hard, just requires you to know what youre doing and not doing, just as importantly.

OpenStudy (lukecrayonz):

So k=5 g(x)=x-6/x+3

OpenStudy (lukecrayonz):

Oh god please tell me I don't have to use the quotient rule.

OpenStudy (psymon):

Not really. Here is how I like to tell people to think of it:  f is still going to be x-6, g being x -3. ANd you donthave to use quotient rule, you can do product rule.